• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 15:05:41 on September 28, 2010

    Entry number 333035

    keyword=Experiment status update

    Target leak has been fixed, we will be doing a hot checkout of the beamline (no beam, just magnet and instrumentation checks) starting at 6pm. We will return to Restricted Access when MCC is done (2--3 hours).

    The CHL is recovering, and the ESR is continuing to cool. So far, so good. Assuming the accelerator comes back online during Wednesday DAY shift, there is a good chance we will lock up and take some beam during Wed SWING / Thurs OWL. The Hall will return to Restricted access during DAY shift Thursday. These plans are very preliminary though, so keep tuned.

    The Hall A RC report with additional detail may be found here: https://hallaweb.jlab.org/wiki/images/a/aa/DVCS-HallA-meeting_28Sep2010.pdf